Two equal mass objects, both traveling at the same speed v, collide as shown to the right. After the collision, the first object travels again with the same speed v, but at a right angle to its initial path. Find the magnitude and direction of the second object’s velocity.
Answer #1

Assuming the objects are travelling towards each other in a straight line. Let the motion of the first object be along positive x axis. Let m be the mass of each of the objects. Let vx be the final speed of second object in x direction and vy be the final speed of second object in y direction.

By conservation of momentum in x direction:

mv - mv = 0+ mvx => vx= 0

By conservation of momentum in y direction:

0 + 0 = mv + mvy => vy = -v

Hence, the second object travels with the same speed v but in the opposite direction to the final velocity of first object.